OpenStudy (anonymous):

Simplify each of the following expressions 5. (-10)+25+(-8) 6.-30+14-8 7. (16)(-9)(3)

OpenStudy (mathlegend):

5. 7

OpenStudy (mathlegend):

6. 8

OpenStudy (anonymous):

can you also show some support with your answer please?

OpenStudy (mathlegend):

Sure

OpenStudy (anonymous):

Thanks :)

OpenStudy (mathlegend):

5. Remember that Order of Operations says that when it is just addition and subtraction you go from left to right

OpenStudy (mathlegend):

-10 + 25 = 15 15 - 8 = 7

OpenStudy (mathlegend):

6. Order of Operations just addition and subtraction again so we move from left to right again

OpenStudy (mathlegend):

-30 + 14 = 16 16 - 8 = 8

OpenStudy (mathlegend):

16 * - 9 = -144 -144 * 3 = -432

OpenStudy (anonymous):

Thanks :)

OpenStudy (mathlegend):

No problem

OpenStudy (anonymous):

can you help with a couple more?

OpenStudy (mathlegend):

Okay sure

OpenStudy (anonymous):

Simpliofy each of the following expressions show your support/work -12.8+2.6+(-11.9)+6.2+0.9

OpenStudy (mathlegend):

So we have addition and subtraction which calls for us to move from left to right

OpenStudy (mathlegend):

-12.8 + 2.6 = -10.2 -10.2 - 11.9 = -22.1 -22.1 + 6.2 = -15.9 -15.9 + 0.9 = -15

OpenStudy (anonymous):

Thanks a bunch :) theres more so ill just hit you up after i copy this down

OpenStudy (mathlegend):

ok

OpenStudy (anonymous):

check outthe new one i just put up